Support activity cost allocation

Jake’s Gems mines and produces diamonds, rubies, and other gems. The gems are produced by way of the Mining and Cutting activities. These production activities are supported by the Maintenance and Security activities. Security costs are allocated to the production activities based on asset value. Maintenance costs are normally allocated based on machine hours. However, Maintenance costs typically correlate more with the number of service calls. Information regarding the activities is provided in the following table:

Maintenance Security Mining Cutting
Number of service calls 17       20    60    20
Machine hours 89       88    182    176
Asset value $200,000       $80,000    $300,000    $300,000
Department cost $25,000       $42,500    $160,000    $95,000

1. Is the process of allocating maintenance costs based on machine hours correct? Identify the measure than can possibly be used to allocate the maintenance costs.

No

Measures that can possibly be used to allocate the maintenance costs.

a. Number of service calls

b. Asset value

c. Department cost

d. None of the above

a

2. Based on your response to part (1), determine the total costs allocated from each support activity to the other activities using the reciprocal services method and the most appropriate cost driver for Maintenance.

Maintenance
Department
Security
Department
Mining
Department
Cutting
Department
Maintenance cost allocation $ $ $
Security cost allocation $ $ $
Production activities total costs $ $

3. Assume Jake’s Gems is considering cutting costs by switching to a simpler support activity cost allocation method. Determine the total costs allocated from each support activity to the other activities.

Particulars Mining
Department
Cutting
Department
Maintenance cost $ $
Security cost $ $
Production activities total cost $ $

Is it viable for the company to choose the direct method of cost allocation?

Which of the following supports your answer?

a. The direct cost allocation method is too simple and it is not accurate.

b. The difference between the direct cost allocation method and the reciprocal method is minimal and the direct method is easier.

c. The reciprocal method of cost allocation is more accurate and easier than the direct method.

d. None of the above.
Answer #1

Answer:

1.jake's gems mines had no the process of allocation maintenance cost on the basis of machine hours is incorrect.Because it is mentioned that maintenance cost correlates to service calls so it should be allocated on the basis of service calls.

2.

Reciprocal Allocation Method Maintenance Security Mining Cutting Total
Number of service calls 20.00 60.00 20.00 100.00
% of number of service calls 20.00% 60.00% 20.00% 100%
Asset value 200,000.00 3,00,000.00 3,00,000.00 800,000.00
% of asset value 25.00% 37.50% 37.50% 100%
Department cost 25000.00 42500.00 160000.00 95000.00 322500.00
let overhead cost of maintenance department is $ x and overhead cost of security is $ y
so, x = 25000+25% of y Equation 1
y= 42500+20% of x Equation 2
or, x =25000+.25y
y=42500+.20x

Working notes:

Substituting the value of x in equation 2:

y=42500+0.20(25000+0.25y)

y=42500+5000+.05y

y-0.05y=42500+5000

0.95y=47500

y=47500/.95

y=$50000

Working note:

substituting the value of y in equation 1:

so, x=25000+25% of y

x=25000+25% of 50000

x=25000+12500

x=$37500

Allocation summary:

Reciprocal Allocation Method Maintenance Security Mining Cutting Total
Number of service calls 20.00 60.00 20.00 100.00
% of number of service calls 20.00% 60.00% 20.00% 100%
Asset value 200000.00 300000.00 300000.00 800000.00
% of asset value 25.00% 37.50% 37.50% 100%
Department cost 25000.00 42500.00 160000.00 95000.00 322500.00
Allocation
Maintenance (on the basis of % of service calls) (37500.00) 7500.00 22500.00 7500.00
Security (on the basis of % of asset value) 12500.00 (50000.00) 18750.00 18750.00
Production activities total cost 201250.00 121250.00 322500.00

3.

Allocation summary:

Direct Method Maintenance Security Mining Cutting Total
Number of service calls 60.00 20.00 80.00
% of number of service calls 75.00% 25.00% 100%
Asset value 300000.00 300000.00 600000.00
% of asset value 50.00% 50.00% 100%
Department cost 25000.00 42500.00 160000.00 95000.00 322500.00
Allocation
Maintenance (on the basis of % of service calls) (25000.00) 18750.00 6250.00
Security (on the basis of % of asset value) (42500.00) 21250.00 21250.00
Production activities total cost 200000.00 122500.00 322500.00

Yes it is viable to choose Direct method of cost allocation.

Because option b .The difference between Direct cost allocation method and Reciprocal method is minimal and the direct method is easier.